OpenStudy (hitaro9):

Let f(x)= x^3 ln(1+x^2) and let the sum from 0 to infinity of anx^n be the taylor series of f about 0. Find a7

OpenStudy (hitaro9):

Eh. I'm just going to give up on this homework. if you want a medal I can give you it

ganeshie8 (ganeshie8):

lol this is easy, u knw maclaurin series ?

OpenStudy (hitaro9):

Not really. But its okay, our teacher lets us drop 1 homework, I'm gunna use it on this one, cause this is the simplest part of it and I'm already totally lost.

OpenStudy (hitaro9):

Thanks anyways, you have a nice day.

ganeshie8 (ganeshie8):

hold up

ganeshie8 (ganeshie8):

taylor series about 0 = maclaurin series

ganeshie8 (ganeshie8):

Maclaurin series : \[f(x) = f(0) + f'(0)x + \dfrac{f''(0)}{2!}x^2 + \dfrac{f'''(3)}{3!}x^3 + ... \]

ganeshie8 (ganeshie8):

simply find f(0), f'(0), f''(0) and plugin ?

ganeshie8 (ganeshie8):

^^ thats the 7th term

zepdrix (zepdrix):

I'll show you the long tedious way to do it, as I showed in the last problem. It's not necessary to take this path, but it's a lot of fun :) check it out!

zepdrix (zepdrix):

\[\Large\rm f(x)=x^3\color{orangered}{\ln(1+x^2)}\] Let, \(\Large\rm \color{orangered}{g(x)=\ln(1+x^2)}\)We want to write that as a power series, then multiply through by the x^3 later on. \[\Large\rm g'(x)=\frac{2x}{1+x^2}\]Yes? So that implies,\[\Large\rm g(x)=\int\limits \frac{2x}{1+x^2}dx\]Again let's move some junk out of the way, get it into geometric form,\[\Large\rm g(x)=2\int\limits x\cdot \color{royalblue}{\frac{1}{1-(-x^2)}}dx\]The blue part is the fun part, we'll rewrite it as a geometric series,\[\Large\rm g(x)=2\int\limits x\cdot \color{royalblue}{\sum_{n=0}^{\infty}(-x^2)^n}dx\]Separating things a bit, and bringing the x into the sum,\[\Large\rm g(x)=2\int\limits \cdot\sum_{n=0}^{\infty}(-1)^n~ x^{2n+1}dx\]

zepdrix (zepdrix):

Bring the integral into the sum,\[\Large\rm g(x)=2 \sum_{n=0}^{\infty}(-1)^n \int\limits x^{2n+1}dx\]Integrating,\[\Large\rm g(x)=2 \sum_{n=0}^{\infty}(-1)^n \frac{1}{2n+2}x^{2n+2}\]

zepdrix (zepdrix):

Recall that this is only g(x), we need to go back to f(x) at this point,\[\Large\rm f(x)=x^3\cdot\color{orangered}{g(x)}\]

zepdrix (zepdrix):

So we bring the x^3 into the sum now, let's divide the 2's out as well,\[\Large\rm f(x)= \sum_{n=0}^{\infty} (-1)^n \frac{1}{n+1}x^{2n+5}\]Phew!

zepdrix (zepdrix):

\[\Large\rm n-th~ term: T_n\]The seventh term is represented by n=6 (since we're starting from 0), \[\Large\rm T_6=(-1)^6 \frac{1}{6+1} x^{2(6)+5}\]\[\Large\rm T_6=\frac{1}{7}x^{17}\]Which works out to the correct values! yay! :)

zepdrix (zepdrix):

A long and tedious approach, but very fun!

zepdrix (zepdrix):

Blah that was sloppy the way I said that, I should have said,\[\Large\rm (n+1)^{th}~term: ~T_n\]

ganeshie8 (ganeshie8):

nice :)